A city government has approved the construction of an $800 million sports arena. Up to $480 million will be raised by selling bonds that pay simple interest at a rate of 5% annually. The remaining amount (up to $640 million) will be obtained by borrowing money from an insurance company at a simple interest rate of 2%. Determine whether the arena can be financed so that the annual interest is $22 million.
Answer by richwmiller(17219)

x+y=800,
0.05*x+0.02*y=22
x=800-y
0.05*(800-y)+0.02*y=22
40-0.05y+0.02*y=22
-0.03*y=-18
y=600 at 2%
x=800-y
x=200 at 5%
x=200 at 5% y=600 at 2%
Yes, it can be done within the limits of 5% <=480 million and 2% <=640
check
0.05*200+0.02*600=22
10+12=22
22=22
ok